OpenStudy (anonymous):

how to simplify sqrt of 18/4

OpenStudy (anonymous):

Divide by the gcd to get \[ \sqrt {\frac{9}{2}} \] then \[ \frac{3}{\sqrt {2}} \]

OpenStudy (anonymous):

* \[\pm \]

OpenStudy (anonymous):

\[3\sqrt{2}/2\]

OpenStudy (turingtest):

yes, if you want to "rationalize the denominator".

OpenStudy (anonymous):

how it comes?

OpenStudy (turingtest):

you mean how do you get from\[3/\sqrt{2}\rightarrow3\sqrt{2}/2 \]? Like this:\[(3/\sqrt{2})(\sqrt{2}/\sqrt{2})=3\sqrt{2}/2\]